OpenStudy (anonymous):

in a right triangle, find the length of the side not given. b=1,c=squrt{17} please help

OpenStudy (anonymous):

a^2 + b^2 = c^2 so, a^2 = c^2 - b^2 = 17 - 1 = 16 a = sqrt (16) = 4